# Understanding Action Effects through Instrumental Empowerment in Multi-Agent Reinforcement Learning

### v1

@@ 5.3 Competition Game
Our approach is flexible to measure the effects of specific groups. We define a fixed agent order $\sigma=(\boldsymbol{p}_{o},\boldsymbol{p}_{t},p_{s})$, where $\boldsymbol{p}_{o}$ are opponent players, $\boldsymbol{p}_{t}$ are teammates, and $p_{s}$ is the striker or the ball player (when the team has possession). Groups $\boldsymbol{p}_{o}$ and $\boldsymbol{p}_{t}$ are processed jointly, and ICV computation is restricted to every second joint action ($\Delta t=2$, see Fig. 2). Since $\pi_{tz}$ is shared across agents, we measure $\bar{\nu}_{d}=\mathrm{JSD}(\pi_{tz}(\cdot\mid s_{t,(k+1)})\,\|\,\pi_{tz}(\cdot\mid s_{t,(k)}))$ as a proxy for strategy change, alongside entropy $\mathrm{H}$ to capture decision uncertainty. The contributions over virtual game time $\tilde{t}$ (in minutes) with seed=175 are shown in Fig. 10. We highlight three key regions, where the situation can be reconstructed to investigate the cause as GRF can store full history. In the gray area, the team loses and regains possession near the penalty area, causing uncertainties and strategy adaption. In the orange region, the opponents scored a goal as reflected by a drop in $\hat{\Phi}(\nu_{v})$ and a rise in $\hat{\Phi}(\bar{\nu}_{d})$, indicating a strong strategy adaptation. In the green region, opponents exert a negative impact on the ball player’s value and determinism just before it scored a goal. In contrast, teammates contribute positively by increasing the ball player’s decisiveness to shoot. While $\hat{\Phi}(\nu_{v})$ and $\hat{\Phi}(\nu_{p})$ contributions do not strictly match, $\hat{\Phi}(\nu_{p})$ and $\hat{\Phi}(\bar{\nu_{d}})$ may still provide meaningful information on the contributions of the players.
@@
(a) Effect of opponents on ball player or striker.
@@
(b) Effect of team players on ball player or striker.
@@
Given the long GRF time horizon ($T=3000$), we compute ICV values over $M=50$ episodes and normalize them by $\kappa=0.04$. We analyze the contributions of opponents (opp) and teammates to the attacking trio, i.e., right midfield (rm), center forward (cf), and left midfield (lf), as shown in Fig. 11(a), and to the ball player (bp) only in Fig. 11(b). We decompose the peak contributions $\hat{\Phi}(\nu_{p})$ to the trio by $\mathrm{H}^{rm}$, $\mathrm{H}^{cf}$, and $\mathrm{H}^{lm}$. We observe that credit got assigned to contributions increasing or decreasing decision certainty that again correlates to the effects on the value. Interestingly, neither opponents nor co-players had a large impact on the ball player on average, making it robust to disturbances by frequently adapting its strategy, as indicated by high $\hat{\Phi}(\bar{\nu}_{d})$.
$\hat{\Phi}(\nu_{v})$   $\mathrm{H}^{rm}$   $\mathrm{H}^{cf}$   $\mathrm{H}^{lm}$
@@ 6 Conclusions
This paper introduced Intended Cooperation Values (ICVs), a novel causal-action attribution method that explains agent behavior through policy influence rather than reward-based evaluations. ICVs quantify how individual actions affect teammates’ decision certainty and preference alignment, effectively capturing implicit cooperation and consensus. The approach proved effective across a variety of environments (cooperative, competitive, and mixed-motive). A major advantage of ICV lies in its simplicity and generality to operate without requiring reward access or architectural changes and can be applied both online and offline. This makes them well-suited for scenarios where interpretability is critical but value functions are unavailable or unreliable. Furthermore, ICVs offer a principled way to analyze intention and influence, grounded in game-theoretic and information-theoretic principles. However, the method also has limitations. Computing ICVs in complex, large-scale environments requires considerable offline post-processing due to the need for sequential action decomposition and intermediate state reconstruction. In addition, our method assumed full state observability, which may limit applicability in some settings. Nonetheless, our method is conceptually extendable to partially observable environments.
@@
Appendix A Ethics Statement.
Our method does not incorporate any form of adversarial attack and pose no risk to human safety. All experiments were conducted in a simulated environment, without involving real-world subjects or raising ethical or fairness concerns. The method is intended to enhance transparency and safety in MARL systems, especially in critical applications.
Appendix B Shapley Value Axioms
@@
Appendix C Proofs
C.1 Advantage
@@
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{v}$ measures the sampled advantage for players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\Lambda^{j}(s_{t,(k)},a_{t}^{i}).$
@@
Given $Q^{j}(s,\boldsymbol{a})=\mathbb{E}_{\boldsymbol{\pi},P}[\sum_{t=0}^{\infty}\gamma^{t}R^{j}\mid S_{0}=s,\boldsymbol{A}_{0}=\boldsymbol{a}]$, the advantage for each agent $j$ is:
$\boldsymbol{\Lambda}^{j}(s_{t,(k)},\boldsymbol{a}_{t})=Q^{j}(s_{t,(k)},\boldsymbol{a}_{t})-V^{j}(s_{t,(k)})=\mathbb{E}_{s_{t+1}\sim P}[R^{j}+\gamma V^{j}(s_{t+1})]-V^{j}(s_{t})$
by noting that our SVMG framework operates between intermediate states $s_{t,(k)}$ at sub-steps $k$, where the time step $t$ is not increased, and noting that rewards are provided only at full next time steps $t+1$, we may omit $R^{j}$ and $\gamma$. Additionally, according to Def. 6 and 8, we consider concrete next intermediate states after agent $i$’s action $a_{t}^{i}$ got processed. We therefore obtain the following advantage for each agent $j$ based on agent $i$’s action (or state propagation):
$\Lambda^{j}(s_{t,(k)},a_{t}^{i})=V^{j}(s_{t,(k+1)})-V^{j}(s_{t,(k)}).$
Then, in the case of centralized critics, we obtain the same result through the Advantage Decomposition Theorem [18], being applicable for all agents at sub-step $k$ for simplicity we use an ordered permutation $k=\sigma(k)$:
$\boldsymbol{\Lambda}^{1:n}(s_{t,(k)},\boldsymbol{a}_{t}^{1:n})$ $=\sum_{k=1}^{n}\Lambda^{k}(s_{t,(k)},\boldsymbol{a}_{t}^{1:k-1},a_{t}^{k})$ (12)
$=\sum_{k=1}^{n}Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})-Q^{1:k-1}(s_{t,(k)},\boldsymbol{a}^{1:k-1})$ (13)
where $Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})=\mathbb{E}_{a^{k+1:n}\sim\pi^{k+1:n}}\left[Q(s_{t,(k)},\boldsymbol{a}^{1:k},\boldsymbol{a}^{k+1:n})\right]$. With the same reasoning as before, $Q^{1:k}(s_{t,(k)},\boldsymbol{a}^{1:k})=V(s_{t,(k+1)})$ and $Q^{1:k-1}(s_{t,(k)},\boldsymbol{a}^{1:k-1})=V(s_{t,(k)})$ yields the advantage at sub-step $k$:
$\Lambda^{k}(s_{t,(k)},a_{t}^{1:k-1},a_{t}^{k})=V(s_{t,(k+1)})-V(s_{t,(k)}).$
C.2 Empowerment
@@
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{p}$ measures the instrumental empowerment as the certainty increase in decision-making of players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)-\mathrm{H}\left(A_{t,(k)}^{j}\right)=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)}).$
@@
We assume $\pi^{j}(\cdot\mid s_{t,(k)},a_{t}^{i})\approx\pi^{j}(\cdot\mid s_{t,(k+1)})$ as otherwise communication-based policies would be required, which are not the scope of this work. At $s_{t,(k-1)}$, by observing action $a_{t}^{i}$ but not executing it to get a new state as per Def. 6, it holds that
$\Delta\nu_{p}$ $=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\log|\mathcal{A}^{j}|-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)},a_{t}^{i}\right)-\log|\mathcal{A}^{j}|+H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)$
$=\sum_{j\in\mathcal{C}_{\sigma}^{i}}H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)},a_{t}^{i}\right)$
$=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{I}\left(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}\right).$ (14)
By setting the probing distribution $p=\pi$, we obtain $\max_{\pi^{i}}\mathrm{I}\left(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}\right)=\mathcal{E}(s_{t,(k-1)})$. ∎

### v2

@@ 5.3 Competition Game
Our approach is flexible to measure the effects of specific groups. We define a fixed order $\sigma=(\boldsymbol{p}_{o},\boldsymbol{p}_{t},p_{s})$, with $\boldsymbol{p}_{o}$ opponents, $\boldsymbol{p}_{t}$ teammates, and $p_{s}$ the striker or ball player. Groups $\boldsymbol{p}_{o}$ and $\boldsymbol{p}_{t}$ are processed jointly, and ICVs are computed at intervals of $\Delta t=2$ (Fig. 2). We measure $\bar{\nu}_{d}=\bar{\mathcal{J}}(\pi_{tz}(\cdot\mid s_{t,(k+1)})\,\|\,\pi_{tz}(\cdot\mid s_{t,(k)}))$ as a proxy for strategy change, alongside $\mathcal{H}$ to capture decision certainty. The contributions over virtual time $\tilde{t}$ $\mathrm{min}]$ with seed=$175$ are shown in Fig. 10. We highlight three regions, where the situation can be reconstructed to investigate the cause. In the gray area, the team loses and regains possession near the penalty area, causing uncertainties and strategy adaption. In the orange region, the opponents scored a goal as reflected by a drop in $\hat{\Phi}(\nu_{v})$ and a rise in $\hat{\Phi}(\bar{\nu}_{d})$, indicating a strong strategy adaptation (scene in Fig. 9(a)). In the green region, opponents exert a negative impact on the ball player’s value and determinism just before it scored a goal (Fig. 9(b)). In contrast, teammates contribute positively by increasing the ball player’s decisiveness to shoot. While $\hat{\Phi}(\nu_{p})$ and $\hat{\Phi}(\bar{\nu_{d}})$ do not strictly match to $\hat{\Phi}(\nu_{v})$, they may still provide meaningful information on players’ contributions.
@@
(a) Contributions of opponents.
@@
(b) Contributions of team players.
@@
Given the long GRF time horizon ($T=3000$), we compute ICV values over $M=50$ episodes, $\Delta t=5$ steps, and normalize them by $\kappa=0.04$. We analyze the contributions of opponents and teammates to the attacking trio, i.e., right midfield (rm), center forward (cf), and left midfield (lf), as shown in Fig. 11(a), and to the ball player only in Fig. 11(b). We decompose the peak contributions $\hat{\Phi}(\nu_{p})$ to the trio by $\mathcal{H}^{rm}$, $\mathcal{H}^{cf}$, and $\mathcal{H}^{lm}$. We observe that credit got assigned to contributions increasing or decreasing decision certainty that again correlates to value effects. The decrease in $\mathcal{H}^{rm}$ by opponents and increase in $\mathcal{H}^{lm}$ by teammates suggest slight insecurities among the right-side players. Interestingly, none of the players exhibited a noticeable average impact on the ball player’s performance $\hat{\Phi}(\nu_{v})$ and decisions $\hat{\Phi}(\nu_{p})$, indicating a robust behavior to disturbances by frequently adapting its strategy, as reflected in high values of $\hat{\Phi}(\bar{\nu}_{d})$.
$\hat{\Phi}(\nu_{v})$   $\mathcal{H}^{rm}$   $\mathcal{H}^{cf}$   $\mathcal{H}^{lm}$
@@ 6 Conclusions
This paper introduced Intended Cooperation Values (ICVs), a novel causal action attribution method that explains agent behavior through policy influence rather than reward-based evaluations. ICVs quantify how individual actions affect teammates’ decision certainty and preference alignment, effectively capturing implicit cooperation and consensus. The approach proved effective across a variety of environments (cooperative, competitive, and mixed-motive). A major advantage of ICV lies in its simplicity and generality to operate without requiring reward access or architectural changes, and it can be applied both online and offline. This makes it well-suited for scenarios where interpretability is critical but value functions are unavailable or unreliable. In such cases, any form of domain knowledge (e.g., being determined to a specific action, as the load action in LBF) may serve as a heuristic to evaluate situations and correctly interpret ICV results. Furthermore, per-step histories of ICVs allow more fine-grained analysis to extract behavioral insights when value feedback is missing, without requiring the situation to be rendered. This is particularly well suited for complex dynamic tasks like GRF. Thus, ICVs offer a principled way to analyze intentions and influences, grounded in game-theoretic and information-theoretic principles. However, the method also has limitations. Computing ICVs in complex, large-scale environments requires considerable post-processing due to the need for sequential action decomposition and intermediate state reconstruction. In addition, our method assumed full state observability, which may limit applicability in some settings. Nonetheless, both SVMG and ICV can still be applied without modification in partially observable domains. However, the resulting ICVs may be less accurate, as some action effects may not be visible to all agents and thus may not alter their policy distributions. More generally, Shapley-based methods inherently involve averaging over all marginal contributions, which can reduce accuracy even in fully observable settings.
{ack}
We gratefully acknowledge the authors of TiZero [18] for providing access to their full model, which was essential for our experiments.
Ethics Statement
Our method is intended to enhance transparency and safety in MARL systems and does not incorporate any form of adversarial attack and pose no risk to human safety. All experiments were conducted in a simulated environment, without raising ethical or fairness concerns.
@@
Appendix A Shapley Value Axioms
@@
Appendix B Propositions and Proofs
In this section, we provide the propositions in a slightly more explicit form, followed by their proofs.
B.1 Advantage
@@
Given the value function $V^{j}(s)$ as defined by Eq. (1) that is learned under $\mathcal{G}_{m}$ dynamics, and employing it as a static predictor within $\mathcal{G}_{s}$, then for each transition of intermediate states $s_{t,(k-1)}\rightarrow s_{t,(k)}$ induced by processing action $a_{t}^{i}$, the value-based marginal contribution, which is given as
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\left[V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})\right],$
quantifies an undiscounted sampled advantage $\hat{\Lambda}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)})$ for players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\hat{\Lambda}^{j}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)}).$
@@
We consider the intermediate states $(s_{t,(k-1)},s_{t,(k)})\in\mathcal{S}$ and the action $a^{i}_{t}\in\mathcal{A}^{i}$ of agent $i=\sigma(k)$, processed according to Eq. (2). The MG value function $V^{j}(\cdot)$ is evaluated statically at concrete state samples, which are either visited or constructed according to Def. 6. This implies no evaluation over all possible future intermediate states. Therefore, we transition from the expectation-based advantage function $\Lambda(s_{t,(k-1)},a^{i}_{t})$ to the sample-based advantage $\hat{\Lambda}(s_{t,(k-1)},a^{i}_{t},s_{t,(k)})$ as in Def. (13). Following the definition of marginal contribution from Def. 8, $\hat{\Lambda}(s_{t,(k-1)},a^{i}_{t},s_{t,(k)})$ is evaluated under the SVMG intermediate-step dynamics for each agent $j\in\mathcal{C}_{\sigma}^{i}$. This implies:
$\sum_{j\in\mathcal{C}_{\sigma}^{i}}\hat{\Lambda}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\left[R^{j}(s_{t,(k-1)},a^{i}_{t},s_{t,(k)})+\gamma V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})\right].$ (12)
As the SVMG framework does not include a reward function $R^{j}$, meaning rewards are only provided at full time-step transitions $t\rightarrow t+1$ and not at intermediate stages $k$, we have $R^{j}(s_{t,(k-1)},a^{i}_{t},s_{t,(k)})=0$ for all sub-step transitions. Thus, Eq. (12) rewrites as:
$\sum_{j\in\mathcal{C}_{\sigma}^{i}}\hat{\Lambda}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\left[\gamma V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})\right].$ (13)
During action processing in the SVMG chain, there is no increment of the time step $t$. Thus, we have $\gamma^{0}=1$, and Eq. (13) becomes:
$\sum_{j\in\mathcal{C}_{\sigma}^{i}}\hat{\Lambda}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\left[V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})\right].$ (14)
By Defs. 8 and 10, the r.h.s. of Eq. (14) corresponds to the value-based marginal contribution $\Delta\nu_{v}$. This proves the proposition. ∎
B.1.1 Remark
The value function $V(s)$ is originally defined for states $s\in\mathcal{S}$ under MG dynamics, where all agents act simultaneously at state $s$. It is learned by the agent’s critic during training as an estimate of the expected return from $s$, as defined by Eq. (1). However, as written in Proposition 1 and shown by Beechey et al. [3], $V(s)$ serves as a static predictor when applying it as feature attribution without ever evaluating the policy $\boldsymbol{\pi}$. Evaluating $V(s_{t,(k)})$ for intermediate states $s_{t,(k)}\in\mathcal{S}$ thus estimates the return as if all agents acted simultaneously at $s_{t,(k)}$ under policy $\boldsymbol{\pi}$ and under MG dynamics. In the SVMG framework, however, the actions of all agents are fixed and do not react to such intermediate updates, as illustrated in Fig. 2. This results in a counterfactual or hypothetical evaluation: if the game were in state $s_{t,(k)}$ and all agents acted according to $\boldsymbol{\pi}$, what would the value be? Based on this reasoning, we leverage the MG-trained value function $V$ within the SVMG framework to quantify agents’ intentional contributions during action processing. The undiscounted sampled advantage $\hat{\Lambda}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)})$, when computed under the SVMG dynamics, thus corresponds to the empirical difference $V(s_{t,(k)})-V(s_{t,(k-1)})$. This approach explicitly serves the purpose of counterfactual attribution.
[…]
However, since our objective is to devise a method that does not require training an additional SVMG-specific model (see Sec. 4), we instead utilize the existing MG-trained value function $V$ directly as a numerical function. This allows us to evaluate $V$ at any encountered or reconstructed intermediate state realization $s_{t,(k)}$, provided $s_{t,(k)}\in\mathcal{S}$, thus making the computation $V(s_{t,(k)})$ both valid and practically feasible.
B.2 Empowerment
@@
For each intermediate state $s_{t,(k)}\in\mathcal{S}$, the entropy-based marginal contribution, given as:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{o}^{i}}\left[\mathcal{H}\left(A_{t,(k+1)}^{j}\right)-\mathcal{H}\left(A_{t,(k)}^{j}\right)\right]\,$,
measures the mutual information $\mathrm{I}(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)})$ as the certainty increase in decision-making of players in $\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{I}(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)})\,.$
Moreover, if the acting policy $\pi^{i}$ maximizes each conditional mutual information, then $\Delta\nu_{p}$ induces the instrumental empowerment:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)},i)\,.$
@@
After plugging $\nu_{p}$ from Def. 11 with peakedness $\mathcal{H}$ into the marginal contribution from Def. 8, we get
$\Delta\nu_{p}$ $=\sum_{j\in\mathcal{C}_{o}^{i}}\left[\mathcal{H}\left(A_{t,(k+1)}^{j}\right)-\mathcal{H}\left(A_{t,(k)}^{j}\right)\right]$ (15)
$=\sum_{j\in\mathcal{C}_{o}^{i}}\left[(\log|\mathcal{A}^{j}|-H\left(A_{t,(k+1)}^{j}\mid s_{t,(k)}\right))-(\log|\mathcal{A}^{j}|-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right))\right]$ (16)
$=\sum_{j\in\mathcal{C}_{o}^{i}}\left[H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)-H\left(A_{t,(k+1)}^{j}\mid s_{t,(k)}\right)\right]\,.$ (17)
We assume that $\pi^{j}(\cdot\mid s_{t,(k)})\approx\pi^{j}(\cdot\mid s_{t,(k-1)},a_{t}^{i})$ holds (as otherwise communication-based policies would be required, which are not the scope of this work.). This implies that the state $s_{t,(k)}$ encodes sufficient information about $a_{t}^{i}$ to enable a consistent measurement of the RV $A^{j}_{t,(k)}$ at sub-step $k$ and state $s_{t,(k-1)}$ by observing the action $a_{t}^{i}$, but without actually executing it to transition to a new state $s_{t,(k)}$, as per Def. 6. Thus, Eq. (17) transforms to:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{o}^{i}}\left[H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)}\right)-H\left(A_{t,(k)}^{j}\mid s_{t,(k-1)},a_{t}^{i}\right)\right]\,$, (18)
which, by the definition of the conditional mutual information, gives:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{o}^{i}}\mathrm{I}\left(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}\right)\,.$ (19)
The instrumental empowerment then trivially follows by Def. 14.